# A local, many-worlds, model of quantum correlations with finite information flow

### v2

@@ Abstract
Ontological theories, such as the de Broglie-Bohm theory, address the measurement problem by introducing auxiliary random variables that specify, in particular, the actual values of macroscopic observables. Such models may be $\psi$-epistemic, meaning the quantum state is not part of the ontology. A serious issue of this route toward a realistic completion of quantum theory is raised by Bell’s proof that ontological theories are nonlocal. A possible resolution is to reject the assumption that measurements have single actual outcomes. Indeed, relaxing this premise, Deutsch and Hayden showed that Bell’s theorem can be evaded by delaying the buildup of the correlations until the parties compare their outcomes at a meeting point. However, the Deutsch-Hayden theory, which is determinist and $\psi$-ontic, leads to an infinite information flow towards the meeting point. Furthermore, alternative branches are weighted by amplitudes, leading to interpretative issues. By integrating the randomness of single-world theories and the branching of the Deutsch-Hayden theory, we introduce a simple $\psi$-epistemic local model of projective measurements on two spatially separate maximally entangled qubits. Because of its randomness, the model requires two ”equally weighted” branches and a finite information flow – just one bit per measurement is communicated to the meeting point. We explore how this hybrid approach, employing both randomness and branching, addresses key challenges of single-world and Deutsch-Hayden theories. On one hand, the branching allows us to circumvent nonlocality and, possibly, contextuality. On the other hand, randomness makes it more natural and economical to derive quantum probabilities from unweighted counts of branches and ensemble averages. Furthermore, it allows for a reduction of the information flow by stripping the quantum state of its ‘ontic’ rank.
@@ I Introduction
The double-slit and Stern-Gerlach experiments are classic textbook examples that illustrate the core puzzle of quantum theory, that is, the superposition and inteference of distinguishable alternative states. According to the Copenhagen interpretation, as long as a system is in a superposition of two or more states, none of them is actualized. Schrödinger famously illustrated the extreme implications of this view with his thought experiment, in which a cat is placed in a superposition of being both dead and alive trimmer . This scenario highlights a fundamental tension between the interpretation and our intuitive expectation that the cat must experience one definite state.
@@
In modern terms, the apparent paradox in Schrödinger’s thought experiment finds a pragmatic resolution in decoherence theory. Since no experiment can detect interference between the ”dead” and ”alive” states, the superposition becomes de facto equivalent to a statistical mixture of the two. This allows for a consistent reduction to one of the two states without leading to contradictions. Thus, it is safe to claim that one of the alternatives is actualized with no future disproval of the claim.
@@
While this approach leads to an interpretation of quantum theory which is practically consistent, it also raises a conceptual problem: how can one state in a superposition be actualized merely on the promise that no future observation will reveal the superposition? It would seem that the system has a kind of foresight of the future. For example, the system could decohere by emitting a photon going into deep space. Since it is plausible that the photon is lost forever, the system could safely ‘decide’ to reduce its state. However how does it know that the photon will not encounter a mirror on one of Jupiter’s moons and be reflected back for reabsorption?
@@
Since the inception of quantum theory, interpretative challenges have raised questions about its completeness epr . To reconcile state superpositions with the definiteness of the our macroscopic experience, a minimal requirement is the existence of auxiliary information that determines the actual macroscopic state that is experienced. This idea is pursued in ‘hidden-variable’ theories and is consistently realized in the de Broglie-Bohm (dBB) theory bohm . In the dBB theory, both the wave-function and the position of the particles specify the actual state of a system. Employing a term coined in the context of quantum foundations, the state represented by the wave-function and the auxiliary variables is referred as an ontic state. More broadly, instead of some variables, the precise ontic state itself could be hidden. Therefore, more generally, we refer to these hypothetical completions as ontological theories.
It is important to emphasize that ontological theories are inherently probabilistic, reflecting the randomness of quantum phenomena. In Bohmian theory, for instance, particle positions are randomly distributed according to Born’s rule. As we will see, however, the randomness of a theory of quantum phenomena is not self-evident.
Specific ontological theories, known as $\psi$-epistemic theories, has gained significant interest over the past decade pbr ; hardy ; leifer0 ; barrett ; leifer ; ring . In this subclass, the quantum state is interpreted as a state of knowledge and is not part of the ontology; instead, its information is encoded in the statistics of the ontic state. This statistical encoding provides a straightforward explanation of why quantum states cannot be cloned or why two non-orthogonal states cannot be discriminated with certainty. Furthermore, there is a link between $\psi$-epistemic theories and classical simulations of quantum communication which employ finite classical communication montina_comm_epist0 ; montina_comm_epist .
@@
However, this route toward a realistic completion of quantum theory leads to a series of conceptual problems, such as contextuality kochen and, more seriously, nonlocality bell , on which we focus in this paper. As shown by Bell, the correlations between two entangled systems are incompatible with any (single-world) local ontological theory; a measurement performed on one system would instantaneously affect the ontic state of the other, in apparent violation of the principle of locality.
Another approach to addressing the interpretative problems of quantum theory is the Everett many-worlds theory everett . Since a cat must be either dead or alive after all, Everett proposed that both states are equally realized, but in separate worlds. This accounts for extremely unlikely yet theoretically possible future interferences between the alternatives. Every time a system enters a superposition of states, it results in a branching into parallel worlds, leading to an exponential blowup of distinct realities over time.
What this theory preserves are the relations between events within each world. However, a fundamental issue arises: where do probabilities come from? Specifically, the theory does not inherently explain why certain events are more probable than others. For example, consider a system evolving into the superposition
$\psi_{+}|+\rangle+\psi_{-}|-\rangle$ (1)

### v3

# Local theories with parallel realities and the epistemic view of the quantum state

@@ Abstract
Hidden-variable theories, such as the de Broglie-Bohm theory, address the measurement problem by introducing auxiliary random variables that specify, in particular, the actual values of macroscopic observables. Such models may be $\psi$-epistemic, meaning the quantum state merely represents a state of knowledge, as opposed to $\psi$-ontic models, in which the quantum state is a physical variable of the theory. A serious issue of this route towards a realistic completion of quantum theory is raised by Bell’s proof that the resulting theories are nonlocal. A possible resolution is to reject the assumption that measurements have single actual outcomes. Indeed, relaxing this premise, Deutsch and Hayden showed that Bell’s theorem can be evaded by delaying the buildup of the correlations until the parties compare their outcomes at a meeting point. However, the Deutsch-Hayden theory, which is deterministic and $\psi$-ontic, leads to an infinite information flow towards the meeting point. Furthermore, alternative branches are weighted by amplitudes, leading to interpretative issues. In this paper, we introduce a general framework that combines the randomness of single-world theories with the coexistence of diverse instances, as found in many-worlds theory. This framework incorporates the existing theories as limiting cases. We explore how this hybrid approach addresses key challenges of single-world and Deutsch-Hayden theories. On one hand, the multiplicity of coexisting instances allows us to circumvent nonlocality and, possibly, contextuality. On the other hand, randomness makes it possible to derive quantum probabilities from unweighted counts of instances and ensemble averages. Furthermore, it can lead to a reduction of the information flow by stripping the quantum state of its ‘ontic’ rank. We illustrate this framework with a simple $\psi$-epistemic local model of projective measurements on two spatially separate maximally entangled qubits. Because of its randomness, the model requires two unweighted instances and a finite information flow – just one bit per measurement is communicated to the meeting point. Setting aside its foundational motivations, this framework has also relevance in quantum communication complexity and leads to novel technical questions, potentially providing new insights into some peculiarities of entanglement.
@@ I Introduction
The double-slit and Stern-Gerlach experiments are classic textbook examples that illustrate the core puzzle of quantum theory, that is, the superposition and interference of distinguishable alternative states. According to the Copenhagen interpretation, as long as a system is in a superposition of two or more states, none of them is actualized. Schrödinger famously illustrated the extreme implications of this view with his thought experiment, in which a cat is placed in a superposition of being both dead and alive trimmer . This scenario highlights a fundamental tension between the interpretation and our intuitive expectation that the cat must experience one definite state.
@@
In modern terms, the apparent paradox in Schrödinger’s thought experiment finds a pragmatic resolution in decoherence theory. Since no experiment can detect interference between the ”dead” and ”alive” states, the superposition becomes de facto equivalent to a statistical mixture of the two. This allows for a consistent reduction to one of the two states without leading to contradictions. Thus, it is safe to claim that one of the alternatives is realized with no future disproof of the claim.
@@
While this approach leads to an interpretation of quantum theory which is practically consistent, it also raises a conceptual problem: how can one state in a superposition be actualized merely on the promise that no future observation will reveal the superposition? It would seem that the system has a kind of foresight of the future. For example, the system could decohere by emitting a photon going into deep space. Since it is plausible that the photon is lost forever, the system could safely ‘decide’ to reduce its state. However how does it know that the photon will not encounter a mirror on one of Jupiter’s moons and be reflected back for re-absorption?
@@
Since the inception of quantum theory, interpretative challenges have raised questions about its completeness epr . To reconcile state superpositions with the definiteness of our unique macroscopic experience, a minimal requirement is the existence of auxiliary information that determines the actual macroscopic state that is experienced. This idea is pursued in “hidden-variable” theories and is consistently realized in the de Broglie-Bohm (dBB) theory bohm . In the dBB theory, both the wave function and the position of the particles specify the physical state of a system. Employing a term coined in the context of quantum foundations, the state represented by the wave function and the auxiliary variables is referred to as an ontic state, since it represents the ontology of the theory. Generally, we refer to these hypothetical completions as ontological theories. In particular, theories in which the wave function is considered a physically real object are called $\psi$-ontic. See Ref. harrigan for a comprehensive discussion on this classification. It is important to emphasize that ontological theories are inherently probabilistic, reflecting the randomness and unpredictability of quantum phenomena. In Bohmian theory, for instance, particle positions are randomly distributed according to Born’s rule.
@@
However, this route towards a realistic completion of quantum theory leads to a series of conceptual problems, such as contextuality kochen and, more seriously, nonlocality bell , on which we focus in this paper. As shown by Bell, the correlations between two entangled systems are incompatible with any single-world local ontological theory; a measurement performed on one system would instantaneously affect the ontic state of the other, in apparent violation of the principle of locality.
The assumption of a single actual outcome of a measurement is an essential hypothesis of Bell’s theorem. Indeed, Deutsch and Hayden (DH) developed a local many-worlds theory deutsch based on Everett’s interpretation everett . In the latter, the wave function is identified as the complete ontology of quantum theory, so that its branching into alternative macroscopically distinct states is a physically real process. Thus, while single-world ontological theories are probabilistic, the many-worlds theory is $\psi$-ontic and deterministic. Randomness arises from the subjective perspective of an observer following a single branch of the tree. By employing the Heisenberg picture, Deutsch and Hayden introduced additional ontological objects – called descriptors – in their local theory. Their argument for locality is clearly illustrated through a simple local realistic model of a ‘nonlocal’ Popescu-Rohrlich (PR) box brassard . The core idea of the argument is to delay the buildup of correlations until the parties compare their outcomes at a meeting point. This delay is possible only if multiple realities coexist locally.
However, since DH theory is $\psi$-ontic, there is an infinite flow of information from the parties to the meeting point. Furthermore, due to determinism, the theory faces the same issue as Everett’s: where do quantum probabilities come from? Although there is extensive literature on the derivation of Born’s rule, no broad consensus exists on how quantum probabilities emerge in Everett’s interpretation. Without entering into this debate, we simply assume – within the context of the many-worlds theory – that branches are weighted by an amplitude, and our expectation of being in a particular branch is given by Born’s rule, which we adopt as a postulate. Occasionally, we will also consider the branch counting proposed by Saunders saunders , since this has some similarity with the counting that we will employ in the framework discussed in this work. Motivated by both foundational considerations and communication-complexity questions, our aim is to investigate alternative local theories that employ fewer resources and unweighted counting rules. To this end, we introduce a general framework that encompasses both single-world and many-worlds theories as limiting cases.
In the search for an alternative local theory with multiple instances, it is fundamental to emphasize that the branching of the wave function constitutes a physical process in both the dBB and Everett theories. In the dBB theory, however, the positions of the particles determine which branch is actually experienced. For this reason, some proponents of the many-worlds interpretation argue that dBB is effectively a many-worlds theory augmented with additional (and, in their view, unnecessary) variables. See, for example, the discussion in Ref. brown and references therein. Indeed, it may seem peculiar that alternative branches of the wave function – each containing detailed biophysical processes, such as brain activity – exist physically, yet are empty of observers. These brain activities in “empty” branches would be akin to “philosophical zombies” introduced in philosophy of mind and in debates about consciousness.
With this important premise on the dBB theory, one might wonder whether any completion of quantum theory is essentially just a many-worlds theory augmented with extra-variables. The crucial point of the question lies in the interpretation of the wave function. If the wave function is regarded as part of the ontology, then it makes sense to claim that the various branches of the wave function correspond to physically existing realities, even if observers, attached to the extra-variables, may follow only one of them. Conversely, each branch of the wave function does not necessarily have physical reality in specific ontological theories, known as $\psi$-epistemic theories, which have gained significant interest over the past decade pbr ; hardy ; leifer0 ; barrett ; leifer ; ring . In this subclass, the quantum state is interpreted as a state of knowledge and is not part of the ontology; instead, its information is encoded in the statistics of the ontic state. Because of this encoding, $\psi$-epistemic theories are essentially equivalent to classical simulations of quantum communication which employ finite classical communication montina_comm_epist0 ; montina_comm_epist . In a $\psi$-epistemic ontological theory, the coexistence of multiple realities would not be justified by the branching of the wave function, but rather by Bell’s theorem. Furthermore, in such theories, the branching of the wave function would not necessarily imply a branching into distinct realities. Since $\psi$-epistemic ontological theories are the only alternative to $\psi$-ontic many-worlds theories with possible extra-variables, we focus on them. An exception worth noting is the many interacting worlds (MIW) approach of Ref. hall , which is $\psi$-ontic in some sense, as discussed later. However, it is not equivalent to many-worlds theory, since the equivalence to quantum mechanics holds only in the limit of infinite worlds (an interesting question regarding this approach will be discussed in the conclusions).
In this paper, we explore whether integrating the randomness inherent in single-world ontological theories with the coexistence of multiple instances – as in the many-worlds theory – can yield a more general framework, while preserving the locality of the Deutsch–Hayden (DH) theory. In particular, the introduction of randomness is a prerequisite for a finite information flow, which is infinite in the DH theory. Moreover, relaxing the determinism of the many-worlds theory makes it possible to derive quantum probabilities through a simple unweighted count of instances and an ensemble average. We illustrate this hybrid framework that incorporates both randomness and multiplicity of instances by showing that just one classical bit of information about the chosen measurement and two unweighted instances are sufficient to simulate local projective measurements on maximally entangled qubits. This model raises the question whether a local $\psi$-epistemic theory can be built.
Our model directly builds on the Toner-Bacon model, which simulates maximally entangled states using one bit of communication and shared randomness toner . By selecting an instance at random, our model reproduces the same statistics as the single-world model in Ref. gisin , in which outcomes are generated through shared randomness and the single use of a nonlocal PR box.
In Sec. II, we discuss the general framework of a single-world ontological theory. In Sec. III, the framework is extended to the case of multiple coexisting realities. This extension includes many-worlds theory as limiting case. In Sec. IV, a simplified version of DH theory is discussed in the two-party Bell scenario. In Sec. V, we introduce our model with finite information flow. A discussion follows in Sec. VI. Finally, we draw the conclusions and perspectives. In particular, we remark that the introduced framework offers a generalized setting for quantum communication complexity, which has its own appeal beyond foundational motivations.
II Single-world ontological theories
Ontological theories of quantum mechanics have been motivated by the apparent contradiction between the well-defined macroscopic reality and the linear evolution of the wave function, leading to the measurement problem. Since the cat in Schrödinger’s thought experiment must be dead or alive after all, some auxiliary information about his actual state should be added to the information contained in the wave function, provided that the linearity of the Schrödinger equation is assumed universal. This completion is consistently realized, for example, in the dBB theory bohm , in which both the wave function and the position of the particles are part of the ontology. In this respect, the dBB theory could be considered a kind of many-worlds theory brown with extra-variables labeling which branch is experienced. Is any completion of quantum theory just a many-worlds theory augmented with extra-variables? That is, should the wave function be considered a physically real entity in any ontological theory? In the context of quantum foundations, such completions, typically referred to as $\psi$-ontic, are opposed to $\psi$-epistemic theories, where the wave function represents only an observer’s knowledge or information about a system. This distinction is extensively discussed in the seminal paper by N. Harrigan and R. W. Spekkens harrigan , who advocate for the $\psi$-epistemic view. Let us quote the closing lines of their abstract: “[…] Einstein was seeking not just any completion of quantum theory, but one wherein quantum states are solely representative of our knowledge”. In particular, the authors point out that one of Einstein’s arguments for incompleteness of quantum mechanics leads to support the epistemic view of quantum states under the premise of local causality. Unfortunately, Einstein’s argument alone is not supportive of the $\psi$-epistemic view, since local causality contradicts quantum mechanics under the condition of single actual outcomes, as shown by Bell. However, there are other reasons for promoting a $\psi$-epistemic view spekkens . Furthermore, $\psi$-epistemic models have an important role in quantum communication complexity, since they are essentially equivalent to classical simulations of quantum communication that employ finite classical communication montina_comm_epist0 ; montina_comm_epist . Since the $\psi$-epistemic view is the only possible alternative to an ontological theory which is not just many-worlds theory with extra-variables, it merits further consideration.
An ontological theory of quantum phenomena is characterized by an ontological space $\Omega$ of ontic states, denoted by $\lambda$, which are interpreted as physically real states of the system. Generally, ontological models are introduced in the context of a quantum system $S$, which undergoes a preparation and a subsequent measurement harrigan . In that context, the ontic state refers to the quantum system $S$, whereas measurements are described by indicator functions, which are conditional probabilities of getting an outcome given $\lambda$. As our focus is on a theory rather than a model, we include macroscopic systems interacting with $S$ in the description. Thus, the ontic state contains, in particular, full information about macroscopic facts, such as the position of the pointer of a measuring device. For example, in Schrödinger’s thought experiment, there exists a subset $\Omega_{A}\subset\Omega$ of the ontological space such that $\lambda\in\Omega_{A}$ if and only if the cat is alive.
Ontic states are typically assumed to evolve deterministically, or more generally, according to some causal rule, such as a Markov process. Since quantum theory is fundamentally probabilistic, the preparation of a quantum state $\psi$ does not uniquely determine the ontic state $\lambda$, but rather specifies a probability distribution $\rho(\lambda|\psi)$, possibly conditioned on additional contextual parameters, that is, we have the mapping
$\psi\rightarrow\rho(\lambda|\psi).$ (1)
Crucially, the converse is also true in $\psi$-epistemic ontological theories, that is, the ontic state $\lambda$ does not uniquely determine the wave function. This is equivalent to state that the wave function is not an element of the ontology. Within this framework, the branching of the wave function is not necessarily a physical process, since the wave function itself is not taken to represent a physical entity. Instead, its role is epistemic: its information is statistically encoded in the distribution of the underlying ontic state. Suggesting that a wave function branches into multiple realities would, in this view, be akin to claiming that a tossed coin splits into two worlds merely because each outcome is possible. Thus, a $\psi$-epistemic ontological theory is not a many-worlds theory with extra-variables. The coexistence of multiple realities will be an additional hypothesis justified by Bell nonlocality, rather than the branching of the wave function.
III Ontological theories with multiple coexisting realities
Curiously, the term $\psi$-epistemic has traditionally been used only in the context of single-world theories. In contrast, within the many-worlds community, the view that the wave function has ontological significance appears to be ubiquitous. For example, in Ref. kent , the ontological status of $\psi$ is identified as one of the two essential characteristics of the galaxy of many-worlds interpretations. Kent states: “ DeWitt (writes) ‘[…] The symbols of quantum mechanics represent reality just as much as do those of classical mechanics.’ There seems to be no dispute in the literature on this point: if a theory is not mathematically realist then it is not a many-worlds interpretation.” Indeed, the ontology of the wave function motivates the branching into many parallel worlds. As far as we know, the epistemic view has never been considered in a scenario involving multiple coexisting realities. Remarkably, the model presented in Ref. hall replaces the wave function with many interacting worlds, so that the quantum state is not treated as an element of reality in the theory. However, the model – which is nonlocal – reproduces quantum mechanics exactly only in the limit of infinitely many worlds. In this limit, the collective classical state of the infinite worlds contains complete information about the quantum state. Thus, the theory becomes $\psi$-ontic as the discrepancy with quantum predictions vanishes.
In this section, we extend the framework of single-world ontological theories by assuming that multiple realities may coexist. These theories are necessarily probabilistic as long as the number of realities is kept finite. This contrasts with the many-worlds theory, where probabilities must be justified or postulated a posteriori. The determinism of many-worlds theory can be recovered in the limit of infinite realities. As previously said, if the considered framework is $\psi$-epistemic, then the branching is not motivated by the branching of the wave function, rather by Bell’s theorem. Our aim is to explore whether quantum correlations can be understood in a locally causal framework, simply by relaxing the assumption of a single realized world without adopting the DH many-worlds theory as a whole.
This section is organized as follows. In Sec. III.1, we extend the single-world ontological theories by introducing multiple coexisting instances of some macroscopic physical quantity. In Sec. III.2, we clarify the terms instances and branches in our context by introducing an alternative representation using occupation numbers, which is more reminiscent of the formalism of many-worlds theory, in particular of Saunders’ branch counting. Furthermore, we show that many-worlds theory can be recovered (with some extra structure) in the limit of infinite instances. In Sec. III.3, the combination of two macroscopic quantities is discussed, which is preparatory for the definition of separability and locality of the theory, given in Sec. III.4.
III.1 Multiple instances of a macroscopic quantity